# Eclipse timing variations in 7 post-common envelope binaries: an update on recent circumbinary models

### v1

@@ Abstract
Historically it is found, and particularly so for PCEBs, that circumbinary models have been unable to predict eclipse times into the future by more than a few years (Pulley et al. 2018; Pulley et al. 2022) and alternative mechanisms for generating these ETVs have been considered. These include apsidal motion, angular momentum loss through gravitational radiation and/or magnetic breaking and magnetic effects within the secondary star e.g. Applegate (1992) and Lanza (2020).
@@ 4 ORIGINS of ETVs
The observed variations in eclipse times of binary systems can be driven by a variety of astrophysical and dynamical phenomena including (a) stellar winds and magnetic breaking (b) gravitational radiation (c) apsidal motion (d) relativistic effects (e) the presence of circumbinary objects and (f) magnetic activity within the secondary star. These processes broadly fall into two categories: those giving rise to secular period changes and those that yield quasi-cyclical variations in eclipse times that are frequently observed in short period PCEB systems.
@@ 4.1 Secular period changes
Long term continuous period changes can occur through angular momentum loss (AML) resulting in the binary period decreasing which is inferred in the system’s ephemeris by the inclusion of a negative coefficient quadratic term. Stellar winds interacting with the systems magnetic fields can generate magnetic breaking and the consequential AML brings the binary pair closer together. Similarly, as the binary pair orbit each other the emission of gravitational waves (gravitational radiation) carries energy away from the system causing the binary separation and period to decrease.
@@ 4.2.1 Magnetic mechanisms
where is the reduced mass of the binary. Table 3 shows the values of the ratio for the Lanza (2020) model. Again, all the ratios are greater than 1, except for the second modulation in Rattanamala’s 2-planet model for RR Cae (Rattanamala et al., 2023), where . Hence the Lanza mechanism is possibly contributing to the modulation in RR Cae only.
[…]
@@ 4.2.2 Circumbinary companions
As the magnitude of increases the contribution from the two right hand terms become less significant and so for large values of , Eq. 11 approximates to:

### v2

# Eclipse timing variations in 7 post-common envelope binaries: an update and analysis of their recently proposed circumbinary models

@@ Abstract
Historically it is found, and particularly so for PCEBs, that circumbinary models have been unable to predict eclipse times into the future by more than a few years (Pulley et al. 2018; Pulley et al. 2022) and other mechanisms for generating these ETVs have been considered. These alternative mechanisms are many and here we have assumed that the underlying ephemeris can be represented in quadratic form where the quadratic coefficient may be positive, negative or zero. This generalised quadratic ephemeris suggests the presence of secular period changes which are real (or intrinsic) changes within the binary system and most commonly include mechanisms such as (i) mass transfer (ii) mass loss (iii) gravitational effects from the third body and (iv) angular momentum loss (AML) through stellar winds or gravitational radiation. The very short binary period and detached nature of these PCEB systems suggest that mechanisms (i), (ii) and (iii) do not contribute significantly to ETVs in these systems.
Other possible causes of ETVs include quasi-cyclical mechanisms where (i) apsidal motion (ii) LTTE from circumbinary bodies or (iii) magnetic effects within the secondary star e.g. Applegate (1992) and Lanza (2020), are the main contenders. Apsidal motion is considered unlikely in short period PCEBs since the binary orbit is expected to circularise very quickly. An overview of these mechanisms is given in Section 2 of Borkovits et al. (2025) where they include a "spurious" category e.g. ETVs caused by star spots and stellar oscillations. From the perspective of PCEBs, the most commonly referred to mechanisms of ETVs are AML, LTTE from circumbinary companions and magnetic effects.
@@ 4 ORIGINS of ETVs
The observed variations in eclipse times of binary systems can be driven by a variety of astrophysical and dynamical phenomena (see for example Zorotovic & Schreiber (2013), Borkovits et al. (2025) and references therein), including (a) stellar winds and magnetic breaking (b) gravitational radiation (c) apsidal motion (d) relativistic effects (e) the presence of circumbinary objects (f) magnetic activity within the secondary star, (g) abrupt changes due to mass transfer and (h) star spots and pulsations. These processes broadly fall into two categories: those giving rise to secular period changes and those that yield quasi-cyclical variations in eclipse times and frequently observed in short period PCEB systems.
@@ 4.1 Secular period changes
Long term continuous period changes can occur through angular momentum loss (AML) resulting in the binary period decreasing which is inferred in the system’s ephemeris by the inclusion of a negative coefficient quadratic term. Stellar winds interacting with the systems’ magnetic fields can generate magnetic breaking and the consequential AML brings the binary pair closer together. Similarly, as the binary pair orbit each other the emission of gravitational waves (gravitational radiation) carries energy away from the system causing the binary separation and period to decrease. Mass transfer from one star to the other will lead to a period change, although that is unlikely for detached systems such as the binaries studied in this paper. For RR Cae a model developed by Rattanamala et al. (2023) proposed the existence of a cold star spot to fit their data, but as pointed out in Section 3.8, this model does not fit our recent data.
@@ 4.2.1 Magnetic mechanisms
where is the reduced mass of the binary.
Table 3 shows the values of the ratio for the Lanza (2020) model. Again, all the ratios are greater than 1, except for the second modulation in Rattanamala’s 2-planet model for RR Cae (Rattanamala et al., 2023), where . Hence the Lanza mechanism is possibly contributing to the modulation in RR Cae only.
| Binary system | (d) | / | / | / | / | (K) | / | Refs. |
| HS0705+6700 | 0.09564 | 0.75 | 0.483 | 0.134 | 0.186 | 2900 | 0.00219 | 1,2,a |
@@ 4.2.2 Circumbinary companions
As the magnitude of increases the contribution from the two right hand terms become less significant and for large values of , when is greater than the amplitude of the periodic term , Eq. 11 approximates to: